When a fish observe a tree oitside will tree appear taller  shorter or same size?  Give reason. 

Grade:12

1 Answers

Arun
5 years ago

Man will appear taller.

If are still confused then you can use another method: n= Real height/ apparent height

Where, n= refractive index

Therefore, na/nw= H/h

Where, na= refractive index of air, nw=refractive index of water= 3/4, H= real height and, h=apparent height

h=H/¾,

Therefore, h=4/3 H>H…